1 l of air weighs 1.293 g at 0 deg c and at 1 atm pressure. At what temperature 1 L of air at 1 atm pressure will weigh 1 g ?

Grade:10

1 Answers

Bhavya
7 years ago
Dear student
Applying Ideal gas equation in both case and equating them :-
P1V1/ P2V2 = n1RT1/n2RT2
Canceling P,V and R since these are not changing we get,
n1T1 = n2T2
1.293/M * 273 = 1/M *T
M gets cancelled
T = 352.989 K